OpenStudy (anonymous):

Find the value of y. Round your answer to the nearest tenth. cos y 8/13

geerky42 (geerky42):

You mean \(\cos y = \dfrac{8}{13}\)?

OpenStudy (anonymous):

Cos y =8/13 y= cos^-1(8/13) = 52.02